Beta Company purchased machinery on June 1, 2018 for $89,000. It provided the following data: Estimated Residual Value - $5 000 Estimated Service Life – 8 years Estimated Total Work Hours – 42 000 hours Estimated Total Production – 525 000 units Machine Usage (2018) – 5 500 hrs (2019) – 6 000 hrs Machine Production (2018) – 48 000 units (2019) – 55 000 units Beta Company uses a fractional-year policy to the nearest full month and its financial year ends December 31.

Required: a) What is the depreciable base of the machine? [1 mark]

b) Calculate the depreciation expense for the year ending December 31, 2018 and 2019 using the following methods: Round to nearest whole number

i. Straight-line. [3 marks]

ii. Units of output. [3 marks]

iii. Sum of the years’ digits. [5 marks] i

v. Double declining balance. [4 marks]
Answer #1

a) Depreciable Base = 89000-5000 = $84000

b) Calculate following

2018 2019
Straight line 10500*7/12 = 6125 84000/8 = 10500
Unit of output 48000*.16 = 7680 8800
Sum of the year digit 84000*8/36*7/12 = 10889 84000*8/36*5/12+84000*7/36*7/12 = 10305
Double decline 89000*25%*7/12 = 12979 89000*25%*5/12+89000*75%*25%*7/12 = 19005